Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

coin probability problem

Expert replies
by sohailmbaprep » Sun May 27, 2012 5:18 am
Pls explain how to solve this:
q)the probability is 1/2 that a certain coin will turn up heads on any given toss.If the coin is to be tossed three times,what is the probability that on at least one of the tosses the coin will turn up tails ?
Join the discussion
Source: — Problem Solving |

by aneesh.kg » Sun May 27, 2012 5:32 am
sohailmbaprep wrote:Pls explain how to solve this:
q)the probability is 1/2 that a certain coin will turn up heads on any given toss.If the coin is to be tossed three times,what is the probability that on at least one of the tosses the coin will turn up tails ?
P(Atleast One Tail)
= 1 - P(No Tail, All Heads)
= 1 - (1/2)*(1/2)*(1/2)
=[spoiler]7/8[/spoiler]

Alternate/Long Method:
P(Atleast One Tail)
= P(1T,2H) OR P(2T,1H) OR P(3T)
= (3!/2!)/2*2*2 + (3!/2!)/2*2*2 + 1/2*2*2
= (3 + 3 + 1)/8
=[spoiler]7/8[/spoiler]

I know this doubt will arise, so let me try to pre-empt it.
Doubt: Why is there a 3!/2! on the numerator for P(1T,2H)?
Answer: The numerator is the Number of favourable outcomes for getting 1 Tail and 2 Heads. You can have 1 Tail and 2 Heads in the following cases:
T T H
T H T..
etc
The number of such ways = The number of ways of arranging T, T and H = (3!/2!)
Aneesh Bangia
GMAT Math Coach
[email protected]

GMATPad:
Facebook Page: https://www.facebook.com/GMATPad
Join the discussion

by mdavidm_531 » Tue May 29, 2012 9:52 pm
aneesh.kg wrote:
sohailmbaprep wrote:Pls explain how to solve this:
q)the probability is 1/2 that a certain coin will turn up heads on any given toss.If the coin is to be tossed three times,what is the probability that on at least one of the tosses the coin will turn up tails ?
P(Atleast One Tail)
= 1 - P(No Tail, All Heads)
= 1 - (1/2)*(1/2)*(1/2)
=[spoiler]7/8[/spoiler]

Alternate/Long Method:
P(Atleast One Tail)
= P(1T,2H) OR P(2T,1H) OR P(3T)
= (3!/2!)/2*2*2 + (3!/2!)/2*2*2 + 1/2*2*2
= (3 + 3 + 1)/8
=[spoiler]7/8[/spoiler]

I know this doubt will arise, so let me try to pre-empt it.
Doubt: Why is there a 3!/2! on the numerator for P(1T,2H)?
Answer: The numerator is the Number of favourable outcomes for getting 1 Tail and 2 Heads. You can have 1 Tail and 2 Heads in the following cases:
T T H
T H T..
etc
The number of such ways = The number of ways of arranging T, T and H = (3!/2!)
Hi, aneesh,

Can we also use binomial probability distribution with this kind of problem?

If so, how?

Thank you,
Dave
Join the discussion

by eagleeye » Tue May 29, 2012 10:27 pm
I know that you've asked Aneesh but here's the answer. You definitely can. In fact, the way aneesh solved it in the second way is an application of the Binomial distribution.
For p=1/2=0.5 being the probability of a head, the probability of a tail is 1-p=0.5

Then the terms below show 3C0 term for no tails, 3C1 for one tail, and 3C3 for 3 tails.

In the case of 3 tosses, 1 = 3C0*0.5^0*0.5^3 + 3C1*0.5^1*0.5^2 + 3C2*0.5^2*0.5^1 + 3C3*0.5^3*0.5^0

So to find the answer, you could either subtract 3C0*0.5^3 from 1 or add the other 3 terms.

Let me know if this helps :)
Join the discussion

by mdavidm_531 » Tue May 29, 2012 10:29 pm
eagleeye wrote:I know that you've asked Aneesh but here's the answer. You definitely can. In fact, the way aneesh solved it in the second way is an application of the Binomial distribution.
For p=1/2=0.5 being the probability of a head, the probability of a tail is 1-p=0.5

Then the terms below show 3C0 term for no tails, 3C1 for one tail, and 3C3 for 3 tails.

In the case of 3 tosses, 1 = 3C0*0.5^0*0.5^3 + 3C1*0.5^1*0.5^2 + 3C2*0.5^2*0.5^1 + 3C3*0.5^3*0.5^0

So to find the answer, you could either subtract 3C0*0.5^3 from 1 or add the other 3 terms.

Let me know if this helps :)
Thanks, eagleeye!

Sharp as ever!
Join the discussion

by aneesh.kg » Tue May 29, 2012 11:02 pm
eagleeye wrote:I know that you've asked Aneesh but here's the answer. You definitely can. In fact, the way aneesh solved it in the second way is an application of the Binomial distribution.
For p=1/2=0.5 being the probability of a head, the probability of a tail is 1-p=0.5

Then the terms below show 3C0 term for no tails, 3C1 for one tail, and 3C3 for 3 tails.

In the case of 3 tosses, 1 = 3C0*0.5^0*0.5^3 + 3C1*0.5^1*0.5^2 + 3C2*0.5^2*0.5^1 + 3C3*0.5^3*0.5^0

So to find the answer, you could either subtract 3C0*0.5^3 from 1 or add the other 3 terms.

Let me know if this helps :)
To add a little to what eagleeye said,
If there are 'n' number of tosses and 'r' number of these tosses must result in a Head, then
Required Probabilty = (n)C(r)/(2^n) = (n)C(n-r) /(2^n)

where,
nCr = (n)C(n-r) are the coefficients of Binomial Distribution.

But, having said all of that, 'Permutations & Combinations' becomes an interesting topic only when you understand the reason behind using nCr or (n!)/(r!)(n - r)! than just applying the above formula.

Both nCr and (n!)/(r!)(n - r)! are the number of ways of arranging r Heads and (n - r) Tails on the n tosses.
Aneesh Bangia
GMAT Math Coach
[email protected]

GMATPad:
Facebook Page: https://www.facebook.com/GMATPad
Join the discussion

by mdavidm_531 » Tue May 29, 2012 11:56 pm
aneesh.kg wrote:
eagleeye wrote:I know that you've asked Aneesh but here's the answer. You definitely can. In fact, the way aneesh solved it in the second way is an application of the Binomial distribution.
For p=1/2=0.5 being the probability of a head, the probability of a tail is 1-p=0.5

Then the terms below show 3C0 term for no tails, 3C1 for one tail, and 3C3 for 3 tails.

In the case of 3 tosses, 1 = 3C0*0.5^0*0.5^3 + 3C1*0.5^1*0.5^2 + 3C2*0.5^2*0.5^1 + 3C3*0.5^3*0.5^0

So to find the answer, you could either subtract 3C0*0.5^3 from 1 or add the other 3 terms.

Let me know if this helps :)
To add a little to what eagleeye said,
If there are 'n' number of tosses and 'r' number of these tosses must result in a Head, then
Required Probabilty = (n)C(r)/(2^n) = (n)C(n-r) /(2^n)

where,
nCr = (n)C(n-r) are the coefficients of Binomial Distribution.

But, having said all of that, 'Permutations & Combinations' becomes an interesting topic only when you understand the reason behind using nCr or (n!)/(r!)(n - r)! than just applying the above formula.

Both nCr and (n!)/(r!)(n - r)! are the number of ways of arranging r Heads and (n - r) Tails on the n tosses.
Yes, you are right.

That being said, we could only use binomial probability distribution when we have just TWO events("success" and "failure", in this case: heads and tails), right?

But if we have three mutually exclusive events then the binomial probability can't be used. After all, the prefix "bi" means "two"

Ahhh.. this seems like a refresher on probability theory haha!
Join the discussion